CASE STUDY:

Torsion in Solid and Hallow Shaft Application

DIRECTIONS:

Select an application on Torsion Axel Problem, research the problem and submit a report according to

Part 1:

- Present your example of Torsion Application for solid and hallow shaft

- Identify the concepts of the problem

Part 2:

- Calculate each process parameters.

- Select a case possible to your solution

Part 3:

- Calculations
Answer #1

TORSIONAL DEFORMATION OF A CIRCULAR SHAFT Torsion is a moment that twists/deforms a member about its longitudinal axis * By oTorsional Deformation of Circular Bars • Assumptions Plane sections remain plane and perpendicular to the AD torsional axis MTorsional Deformation L 0 = angle of twist B F ө IL R O = shear strain is the shear strain, also remember that tano = 0, thusTorsional Deformation τ Сө 9 RL r Shear stress at any point in the shaft is proportional to the distance of the point from thTorque transmitted by shaft(solid) T I total turning moment due to turning force = total force on the ring x Distance of the0 1 Torque transmitted by shaft(hollow) r total turning moment due to turning force = total force on the ring x Distance of tPower transmitted by shaft Power transmitted by the shafts N = r.p.m of the shaft T = Mean torque transmitted Angular speed oPower transmitted by shaft Power transmitted by the shafts N = r.p.m of the shaft T = Mean torque transmitted Angular speed oTorque in terms of polar moment of inertia T r R Moment dT on the circular ring dr = 250dr xrºx 2.trdr = (dA = 2.rdr) R R xrTorque in terms of polar moment of inertia τ T= - XJ R T T I R J R τ СӨ = RL T τ CO C = Modulus of rigidity J R L e = Angle oPolar Modulus Polar modulus is defined as ration of polar moment of inertia to the radius of the shaft. Z J R л For solid shaTorsional rigidity Torsional rigidity is also called strength of the shaft. It is defined as product of modulus of rigidity (Shaft in combined bending and Torsion stresses м. Г X == JT D Shear stress at any point due to torque T 9 Т Txr =9= J J D SheShaft in combined bending and Torsion stresses Major principal Stress 0 +T 2 32M 2 x D + 32M 2 xD + 167 D 16 D (M+ VM + T2 SOShaft in combined bending and Torsion stresses Major principal Stress 16D M+M+12 [D.* -D Minor principal Stress (M-/M² + T²Application to a Bar Normal Force: Fo Fn Bending Moment: ME M Shear Force: Ft Ft Torque or Twisting Moment: MA G 16 M.a solid bar of circular cross section d 15 in. d=40 mm, L= 1.3 m, G = 80 GPa (a) T= 340 N-m, Tmax 0 = ? (b) Tall = 42 MPa, Pa(b) due to Tall 42 MPa T=rd Tall/ 16 = 1 (0.04 m) x 42 MPa / 16 = 528 N-m due to all 2.5° 2.5 x a rad / 180° 0.04363 rad T 80A hollow shaft dia ratio 3/5 is required to transmit 450Kw at 1200pm, the shearing stress in the shaft must not exceed 60N/mmD min T = 194) 16 2NT P Il 60 P60 T 35.82x10 N/mm ZNTD- TI 35.82x101x16 = TT X 60 3,0427x10 N/mm D-3/5D) 3.0427x10 D p-(1-1296) = 3.0427x10 D 3.0427X10 0972 D 9.106mm